A ball is propelled upward from ground level. After 1 seconds, its height in feet is a function defined by f(t)=-16t^2+64t. After how many seconds will it reach a height of 64 feet?
Answer by nerdybill(7384)

A ball is propelled upward from ground level. After 1 seconds, its height in feet is a function defined by f(t)=-16t^2+64t. After how many seconds will it reach a height of 64 feet?
set f(t) to 64 and solve for t:
f(t)=-16t^2+64t
64=-16t^2+64t
16t^2+64=64t
16t^2-64t+64=0
dividing both sides by 16 we get:
t^2-4t+4=0
(t-2)(t-2) = 0
t = 2 seconds